OpenStudy (anonymous):

A bank account has a beginning balance of $560.00. After 6 months, the balance in the account has increased to $572.60. What interest rate is being earned annually on this account?

OpenStudy (wolf1728):

The compound interest rate formula: http://1728.org/compint2.htm log(1 + rate) = {log(total) -log(Principal)} ÷ Years log(1 + rate) = (log (572.60 -log(560)} ÷ .5 log(1 + rate) = {2.7578513437 - 2.748188027} ÷ .5 log(1 + rate) = 0.0096633167 ÷ .5 log(1 + rate) = 0.0193266334 1 + rate = 10^0.0193266334 = 1.04550625 rate = .04550625 or rate = 4.550625% Calculator to check your work: http://1728.org/compint.htm